# Rethinking Programmed I/O for Fast Devices, Cheap Cores, and Coherent Interconnects

### v2

@@ 6.4. Accelerating Timely Dataflow
Timely Dataflow schedules operators in a user-defined dataflow graph to evaluate complex operations on large streams of data, operating on variable-sized batches. Operators may be arbitrarily complex and are independent, and thus benefit from offloading where the overhead of shipping the data to the FPGA is compensated by greater throughput of the offloaded operators.
@@
To test the limits, we construct a worst-case scenario: a large (31-element) graph of simple operators (filters). The large graph maximizes communication overhead in the form of progress tracking data, while the simple operators leave the FPGA with only a small performance lead over the CPU. We want to predict whether PIO can make FPGA offload work in regimes where the fixed overheads of DMA make it impractical, for example where small batch sizes are needed to ensure data freshness (as in Differential Dataflow (McSherry et al., 2013)).
@@
The first (blue) bars of Figure 11 show the baseline performance of the non-offloaded (CPU-only) system against batch size. Note that the horizontal axis is a log scale; All four curves are actually straight lines, with offsets and slopes obtained by a least-squares fit listed in Table 2 along with the coefficient of determination ($r^{2}$). The remaining bars cover the three offload cases. Each is the median of 1000 samples.
While no offload configuration beats the baseline, the two PIO cases outperform DMA up to batch sizes of 512B (for PCIe) and 4KiB (for ECI), with ECI almost matching the baseline at the smallest batch size (128B). To understand why offloading fails to produce a speedup in this example, and what changes will be necessary to allow it to, we look more closely at the fitted coefficients in Table 2.
As shown by the $r^{2}$ values, the fit for each dataset is excellent. We can thus model the time taken to process a batch as a fixed component, or overhead, independent of the amount of data processed (e.g., running the Timely scheduler, or setting up DMA descriptors) plus a component which scales linearly with the size of the batch.
Compared to the baseline, DMA offload has a significantly lower per-byte cost ($3.7$ vs. $8.7ns/B$) but higher fixed cost ($336\mu s$ vs. $204\mu s$). This includes the cost of descriptor setup and interrupt handling, while the per-byte cost shows the higher throughput of the FPGA. This includes both computation and data transfer but as DMA makes very efficient use of the (PCIe 3.0 x8) interconnect, once setup costs are paid, the per-byte transfer cost is on the order of $0.1ns$, allowing us to estimate the per-byte cost of FPGA compute at $3.6ns/B$. The CPU figure represents only compute.
This linear model predicts a break-even at 26KiB. This matches the observed data (not plotted) where DMA offload first outperforms the baseline at a batch size of 32KiB. Wherever the FPGA compute cost is at least $0.1ns$ lower per byte than the CPU, DMA offload will outperform for large enough batches. However, due to the large fixed overhead it can never outperform the CPU for batches smaller than 16KiB. This limits its usefulness where small batch sizes are needed.
While they make less efficient use of the interconnect bandwidth (as indicated by their higher per-byte costs), their lower fixed overhead allows the PIO modes to significantly outperform DMA at smaller batch sizes. What would it take for the faster of the two (ECI) to outperform the CPU at a moderate batch size?
As the FPGA pipeline is identical in all cases, the limiting factor is the per-byte transfer cost. Taking our above estimate of FPGA compute at $3.6ns/B$, we can estimate this cost at $30-3.6=26.4ns/B$. Letting $S$ stand for batch size and $C$ for this transfer cost, the two break even where
$204\mu s+S\times 8.7ns/B=221\mu s+S\times(C+3.6ns/B)$
Rearranging gives
$C=5.1ns/B-17\mu s/S$
5.1ns/B being the speedup due to FPGA processing and $17\mu s$ the fixed overhead relative the baseline, the negative sign due to the higher-than-baseline overhead.
From this we see that given the overhead the smallest batch for which a speedup is possible is around 3300B, and that a batch size of 4096B demands a transfer cost not more than 0.95ns/B, or 120ns per 128B cache line. As the existing protocol serializes on each line, this cost includes the round-trip latency. As the hardware can handle many in-flight transfers at once, modifying the protocol to transfer 32 lines at once would reduce the per-byte cost enough that ECI-PIO offload breaks even.
Additionally, bringing the fixed cost down reduces the smallest practical batch. If the relative overhead could be brought below 500ns/batch then with parallel cache lines giving a 1ns/B transfer cost, FPGA offload using ECI-PIO would outperform the CPU from a batch size of 128B. With such an optimized protocol, ECI-PIO would outperform all other cases until a batch size of 142KiB, above which PCIe-DMA would overtake it.
Figure 11. Timely Dataflow offload, graph size 31
Table 2. 1st-order (offset + slope) coefficients fitted to the data series of Figure 11, as cost per byte and fixed overhead
Time/Byte (ns) Overhead ($\mu$s) $r^{2}$
Baseline 8.7 204 1.00
PCIe-DMA 3.7 336 0.99
PCIe-PIO 88 277 1.00
ECI-PIO 30 221 1.00
6.5. Offloading Bloom Filters
@@
In a second acceleration experiment we offloaded Bloom filters (Bloom, 1970) to the FPGA. Bloom filters are used to efficiently check whether an element is a member of a set. It uses little memory and allows for fast lookups. A Bloom filter is equipped with $k$ different hash functions and a lookup table. Testing the presence of an element requires computing these $k$ hash functions and querying the lookup table with the results.
@@
We took inspiration from the applications discussed in Fleet (Thomas et al., 2020). As the CPU baseline, we re-implemented the Fleet C++ code in Rust as a Timely Dataflow operator. In the Fleet implementation, the hashes were computed in parallel for each byte sequentially using AVX instructions, which we replicated in Rust using SIMD instructions. On the FPGA side, the hash computation is parallelized and the iteration over the element’s bytes is pipelined. We present the numbers of using a single thread on CPU and a single compute unit on FPGA in order to keep the comparison fair.
With the Bloom filter integrated into Timely Dataflow as a single offloaded operator, most of the additional overhead from processing sub-graphs is eliminated and the flow of data between FPGA and CPU is reduced to the minimal practical batch. In addition, the computational load is significantly higher compared to a simple filter. The $k$ hash functions consist of sequential bit shifting, addition, and XOR-ing, which are very fast in FPGA logic and deliver the offloading benefit comes from. Therefore in this application we concentrate on minimal practical batches and direct offloading benefits.
@@
In Figure 12, the latencies for PIO and CPU both start around 25 $\mu$s, due to the high overhead of streaming the input data. The real processing time of one element is 2.6 $\mu$s on the CPU and 1.7 $\mu$s when offloading with ECI. As overall processing time increases with the data size, we see the gap between CPU and PIO more clearly. After 8 KiB, same as in previous experiments, XDMA starts to perform better than other approaches as performance becomes dominated by the higher bulk throughput of PCIe DMA.
@@ 7.2. Coherent interconnects
Coherent replacements or extensions of existing device interconnects have been under development for some time and are beginning to see widespread availability. IBM’s OpenCAPI (Stuecheli et al., 2018) builds on PCIe by adding a protocol layer for coherence. It and the competing Gen-Z (Keeton, 2015) and CCIX (CCIX Consortium and others, 2024) protocols have meanwhile been either merged into or replaced by CXL (Li et al., 2023), which seems likely to be the standard interoperable protocol in the short- to medium-term future. However, commercial CXL hardware implementing the more recent revisions, which allow less restricted use cases, is still not widely available.
@@ 7.3. Cache line-based communication
Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer where small operations are expected, as in FastForward (Giacomoni et al., 2008) and Barrelfish (Baumann et al., 2009) protocols. More recently Concord (Iyer et al., 2023) communicates scheduling decisions between workers and the dispatcher via a polled cache line, converting worker threads from interrupt-driven to poll-mode “CPU drivers”. A similar technique is used in Shinjuku (Kaffes et al., 2019).
@@ 7.4. Exploring PIO
In kPIO+WC (Larsen and Ben Lee, 2015) the authors demonstrate the benefit of write combining for PCIe-based PIO. Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small and medium-sized messages and comparable throughput for large messages. The ThunderX-1 automatically performs combining for PCIe writes, and the results of Section 6.2 thus reflect this optimization.
@@ 8. Conclusion and Future Work
In this paper we have demonstrated that PIO on modern hardware outperforms DMA over a wide range of payload sizes in a number of different applications. For anything up to a disk block, VM page, or Ethernet Jumbo Frame, PIO offers (often dramatically) lower transfer times mostly thanks to its low fixed overheads. This advantage is consistent across RPC workloads, NIC packet transfer, and FPGA offloaded applications.
This strongly contradicts the accepted wisdom, under which these specific payloads and applications are the canonical examples of DMA-suitable tasks. Moving beyond conventional interconnects (PCIe), we have shown that the advantage grows even greater for the emerging class of coherent interconnects. This is partly due to a more suitably-sized unit of transfer (the cache line) and partly due to the opportunity to interact with the coherence protocol in a carefully-optimized manner.
Moving back to PIO makes sense in a landscape of plentiful, cheap CPU cores and fine-grained RPC-style workloads.
This work is ongoing, and we intend to develop it further in several ways. Both the cache-coherent NIC and Timely Dataflow offload designs will continue to be optimized, to explore the limits of PIO-style device communication on emerging coherent interconnects. We will also further explore the design of appropriate programming interfaces and OS-level integration of coherent mechanisms for high-performance devices.

### v3

@@ 5.3. Offloading Timely Dataflow
Finally, we look at hardware offload of operators in a real-time stream processing system: Timely Dataflow (Murray et al., 2013; McSherry, 2025). Timely schedules operators in a user-defined dataflow graph to evaluate complex functions over on data streams, operating in variable-sized batches. Operators are independent and may be arbitrarily complex, thus benefiting from hardware offloading where the performance improvement outweighs the overhead of shipping data to the accelerator.
Timely also illustrates the flexibility of the ECI PIO protocol: our offloading implementation the protocol in Section 4 synchronously to exchange statistics between FPGA and CPU before and after processing a batch, and operators asynchronously for processing data in the batch in parallel.
We modify the latest Rust implementation of Timely (McSherry, 2025) to include the ability to partition its dataflow graph between software operators and those implemented on the FPGA, with automatic insertion of communication channels between the CPU and FPGA to transfer both data and progress tracking information across the hardware/software boundary. We perform two experiments: an extreme synthetic benchmark with little actual computation, and real operators in the form of offloaded Bloom filters.
Communication-only graph
@@
The synthetic benchmark uses 31-element pipeline graph of trivial filter operators (filters). This large operator graph maximizes communication overhead in the form of Timely progress tracking data, while the simple operators mean the FPGA has little performance advantage (if any) over the CPU, mimicking cases where small batch sizes are need to ensure data freshness, as in Differential Dataflow (McSherry et al., 2013).
Figure 11. Timely synthetic filter graph offload
@@
The first (blue) bars of Figure 11 show the baseline performance of the non-offloaded (CPU-only) system against batch size. The remaining bars cover the three offload cases.
ECI PIO batch latency is lower than both PIO and DMA over PCIe for all batch sizes, even a single cache line (128B). Moreover, PIO over ECI is the only technique that delivers lower latency than the software-only Rust implementation, even in this near-worst-case scenario for hardware offload.
Bloom filters
@@
In our second Timely experiment we offload Bloom filters (Bloom, 1970) to the FPGA, replicating the application scenario used to evaluate Fleet (Thomas et al., 2020). Bloom filters are used to efficiently check whether an element is a member of a set, using little memory. A Bloom filter consists of $k$ different hash functions and a lookup table. Testing the presence of an element requires computing these $k$ hash functions and querying the lookup table with the results.
@@
In the Fleet C++ software-only baseline, the hashes were computed in parallel for each byte sequentially using AVX instructions. We re-implemented this in Rust as a Timely operator using Arm SIMD instructions. On the FPGA side, the hash computations are parallelized and iteration over the element’s bytes is pipelined. We measure a single CPU thread and compare it with a single offloaded operator.
@@
In contrast to the synthetic benchmark above, this scenario has much lower progress tracking overhead (only a single operator is offloaded, compared with an entire sub-graph) and much higher computation cost (the Bloom filters themselves). The $k$ hash functions consist of sequential bit shifting, addition, and XOR-ing, which are fast in FPGA logic and deliver the offloading benefit even for small batches.
@@
Figure 12 shows the results. Latency for both PIO solutions and software (CPU) all start around 25 $\mu$s, due to the high overhead of streaming the input data. The per-element processing time is 2.6 $\mu$s on the CPU and 1.7 $\mu$s when offloading with ECI. As the batch size (and hence processing time) increases, the advantage of hardware offload increases, but so does the cost of PIO over PCIe. DMA latency barely increases, but remains significantly higher than that for ECI.
@@
As with the previous experiments, this shows the advantages of the efficient, low-latency communication between CPU and device that is possible using a cache-coherent interconnect, over and above both PIO and DMA over an interconnect like PCIe. We anticipate that a future implementation of CXL.mem 3.0, which similarly exploited cache coherence at the message level, would deliver similar benefits.
@@ 6.2. Coherent interconnects
Coherent replacements or extensions of existing device interconnects have been under development for some time and are beginning to see widespread availability. IBM’s OpenCAPI (Stuecheli et al., 2018) builds on PCIe by adding a protocol layer for coherence. It and the competing Gen-Z (Keeton, 2015) and CCIX (CCIX Consortium and others, 2024) protocols have meanwhile been either merged into or replaced by CXL (Li et al., 2023), which seems likely to be the standard interoperable protocol in the short- to medium-term future. However, commercial CXL 3.0 hardware implementing the more recent revisions has yet to appear.
@@ 6.3. Cache line-based communication
Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer for small operations, as in FastForward (Giacomoni et al., 2008) and Barrelfish (Baumann et al., 2009) protocols. More recently Concord (Iyer et al., 2023) communicates scheduling decisions between workers and the dispatcher via a polled cache line, converting worker threads from interrupt-driven to poll-mode “CPU drivers”. A similar technique is used in Shinjuku (Kaffes et al., 2019).
@@ 6.4. Exploring PIO
Larsen and Lee (Larsen and Ben Lee, 2015) demonstrate the benefit of write combining for PCIe-based PIO. Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small-medium-sized messages and comparable throughput for large messages. The ThunderX-1 automatically performs combining for PCIe writes, and the results of Section 5.1 thus reflect this optimization. Current work (Liew et al., 2025) proposes a write-reordering unit in the device to simplify the use of fast PCIe posted writes and using Ensō (Sadok et al., 2023) DMA to transfer data from the device.
@@ 7. Conclusion and Future Work
It is important to regularly question long-held assumptions about systems, particularly in the light of new technological trade-offs and emerging workloads.
We show that modern workloads can benefit from techniques that optimize for small, frequent, low-latency interactions between CPU cores and devices as much as the more traditional large, throughput-oriented interactions for which current PCIe interconnects and DMA are optimized.
Moreover, the use of cache coherent interconnects between CPU cores and devices enables new techniques with this property, particularly when the device is able to participate in a standard cache coherence protocol at the level of individual cache transitions rather than naive coherence.
We have demonstrated the potential of these techniques by showing a family of messaging protocols implemented over Enzian’s cache coherent interconnect, but more importantly this opens up a wide design space for the use of future coherent protocols like CXL.mem 3.0, which offers many possibilities for more efficient fine-grained communication within and between machines.